# DynamicKV: Task-Aware Adaptive KV Cache Compression for Long Context LLMs

### v2

@@ 5 Experiments
We conduct comprehensive comparative and ablation experiments to verify the effectiveness of our DynamicKV. In Section 5.1, we introduce the models, datasets and baselines used in our experiments. Section 5.2 provides a performance comparison between DynamicKV and baseline approaches. Next, in Section 5.3, we present the results of DynamicKV on the Needle in Haystack Task. Finally, in Section 5.4, we conduct an ablation study on the parameters of our method to validate its feasibility.
5.1 Implementation details
@@
LongBench is a comprehensive benchmark for evaluating the contextual understanding capabilities of LLMs. For our comparative experiments, we use 16 English datasets from this benchmark, specifically NarrativeQA (Kočiskỳ et al., 2018), Qasper (Dasigi et al., 2021), MultiFieldQA-en, HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), GovReport (Huang et al., 2021), QMSum (Zhong et al., 2021), MultiNews (Fabbri et al., 2019), TREC (Li and Roth, 2002), TriviaQA (Joshi et al., 2017), SAMSum (Gliwa et al., 2019), PassageCount, PassageRetrieval-en, LCC (Guo et al., 2023), and RepoBench-P (Liu et al., 2023). These cover key long context application scenarios such as Single-Document QA, Multi-Document QA, Summarization, Few-shot Learning, Synthetic Tasks, and Code Completion. Additionally, for the experiment on the Needle in Haystack task, we test the models across their maximum length ranges [8k, 32k, 1M] using the PaulGrahamEssays dataset.
@@
We evaluate the recent fixed-pattern token-dropping methods, including: (1) StreamingLLM, which utilizes attention sinks and rolling KV caches to retain the most recent tokens. (2) H2O, which employs a Heavy Hitter Oracle for KV cache eviction. (3) SnapKV, which selects important tokens for each attention head through clustering. (4) PyramidKV, which introduces a pyramid pattern where layers select important tokens in a monotonically decreasing manner.
@@ 5.2 Comparative experiments on LongBench
Model Size Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
@@
Table 1: Performance comparison on the LongBench dataset for full KV cache, previous methods (StreamingLLM, H2O, SnapKV, PyramidKV), and our DynamicKV method, with KV cache sizes of 128 and 512, using models including LLaMA3-8B-Instruct, Mistral-7B-Instruct-v0.2, QWen2-7B-Instruct, and InternLM-2.5-Chat-1M. Bold indicates the best performance.
With the total KV cache size fixed at 128 and 512, we compare the performance retention of StreamingLLM, H2O, SnapKV, PyramidKV, and our proposed method, DynamicKV, relative to FullKV. As shown in Table 1, DynamicKV demonstrates stable improvements even while maintaining an extremely low KV cache size relative to the total context (128: 1.7%; 512: 6.9%). Specifically, with the cache size of 128, DynamicKV outperforms the best alternative by 0.3%, 0.97%, 1.68%, and 0.79% on LLaMA, Mistral, Qwen, and InternLM, respectively, retaining 90%, 87%, 78%, and 83% of the overall performance. Moreover, with a cache size of 512, DynamicKV surpasses the highest-performing method by 0.43%, 0.19%, 0.69%, and 0.53% on the same models, retaining 97%, 96%, 96%, and 89% of FullKV’s performance. The data in the table clearly demonstrate DynamicKV’s effectiveness under extreme compression, achieving nearly FullKV-level performance with just 6.9% of the cache size. The experimental results show that DynamicKV can improve the effect of complex tasks such as code completion more obviously based on maintaining PyramidKV performance, and greatly improve the performance upper limit of lower KV cache size.
@@ 5.3 Visualization on Needle-in-Haystack Task
Figure 4: Performance Comparison on the Needle in a Haystack Task Using Mistral-7B-Instruct-v0.2.
The needle-in-a-haystack test involves inserting key information at random positions within a long context and setting answers to evaluate whether LLMs can accurately detect critical information in extensive contexts. To further illustrate the effectiveness of our approach in compressing the KV cache, we conduct additional experiments using Mistral on the needle-in-a-haystack task, focusing on maintaining an optimal size for the KV cache.
As shown in Figure 4, we insert information at various positions in the Paul Graham Essays dataset and extract answers by prompting the model to generate responses. The green blocks indicate that the response matches the contents of the needle, but the colour change from yellow to red indicates that the response is more irrelevant to the needle.
We test a fixed KV cache size of 64 using FullKV, StreamingLLM, H2O, SnapKV, PyramidKV, and the DynamicKV method. The results indicate that DynamicKV maintains 90% of the model’s performance even under extreme compression, improving accuracy by 57%, 37%, 41%, and 11% compared to the other methods, respectively. Additionally, the figure shows that with a context length of up to 7000, the extreme compression of DynamicKV nearly achieves full scores, and even beyond 7000, it shows significant improvements compared to other approaches. This finding illustrates that DynamicKV has a distinct advantage in hierarchical token selection and confirms that the number of critical tokens contained at different layers is always dynamic.
5.4 Ablation Study
KV size LlaMA-3-8B- Instruct Mistral-7B- Instruct-v0.2 Qwen2-7B- Instruct InternLM2.5-7B- Chat-1M
64 34.93 33.95 32.67 33.67
96 36.70 36.22 34.85 35.31
128 37.75 37.33 35.82 36.07
256 39.83 39.23 36.98 37.29
512 40.73 40.90 39.16 38.39
1024 41.22 41.48 39.72 38.86
Table 2: Performance of DynamicKV with different KV cache size.
In this study, we investigate the performance of the DynamicKV mechanism across varying key-value cache sizes. The results, as shown in Table 2, reveal a consistent improvement in performance with an increase in the cache size for all evaluated models. For the LlaMA-3-8B-Instruct, the performance metric improved from 34.93 to 41.22 as the key-value cache size was increased from 64 to 1024. This improvement is also applicable to other models. These findings underscore the effectiveness of the DynamicKV cache in leveraging KV cache compression to maintain the capabilities of long context. Notably, a larger cache capacity is generally associated with superior performance. Nonetheless, it is essential to strike a balance when selecting the cache size, taking into account the practical constraints related to storage and computational resources.
☞ More analyses in Appendix
In addition to the above discussions, we conduct more related analyses and show them in Appendix: (1) DynamicKV is effectively scalable to other models on the Needle-in-Haystack task (Appendix A.3); (2) Enhancing generation efficiency with 22.5-129% higher token throughput compared to fullKV and superior memory optimization saving up to 11.2% GPU memory while maintaining stable performance (Appendix A.4); (3) Task-Aware adaptability that dynamically adjusts token processing across transformer layers based on input complexity(Appendix A.5). Please refer to the Appendix for more details.
@@ 6 Conclusion
In this study, we analyze the intrinsic patterns exhibited by large language models (LLMs) when processing long-context inputs across different task types. Our empirical findings reveal significant variations in the distribution of attention across these task types. Based on this observation, we introduce DynamicKV, a novel layer-aware KV cache compression approach that dynamically adjusts the KV cache size across layers. We evaluate the effectiveness and generalizability of DynamicKV through experiments on 16 datasets from the LongBench benchmark, demonstrating its broad applicability and performance benefits. From the results, we mainly conclude that: (1) a wave-like pattern is followed in complex reasoning tasks (e.g., code completion tasks); (2) a pyramid-like pattern is followed in Synthetic and Summarization tasks; (3) The dynamic hierarchical adaptive DynamicKV approach is capable of formulating a relatively appropriate KV cache retention strategy in accordance with diverse tasks. Particularly, in the circumstance of maintaining an extremely small KV cache size, the effect is significantly enhanced.; In the future, we hope that there is a more suitable method to perform KV cache compression without increasing the computation.
@@ A.1 Model Details
All the model structures and details in our experiment are shown in Table 3.
@@ A.2 Dataset Details
The data sources, average length, evaluation metrics, language, and data volume of the LongBench(Bai et al., 2023) dataset’s subdatasets are shown in Table 4.
@@
Table 5: Configuration of Models.
@@ A.3 Need in a HayStack
The experimental results are presented in Table 5, which illustrates the performance of our DynamicKV approach across various models, including LlaMA and Qwen, in Needle in a Haystack task.
@@ A.4 Efficiency Experiments
Dataset Layer-wise ( 1 - 32 ) Tokens
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32
NarrativeQA 93 75 111 51 71 75 91 108 97 154 138 143 196 126 156 201
173 133 202 200 158 128 115 86 123 109 117 76 122 120 142 206
Qasper 157 66 92 43 59 70 77 111 93 137 132 160 187 122 147 186
184 132 183 181 160 135 122 77 119 117 128 69 127 113 179 231
MultifieldQA-en 245 79 109 50 76 75 92 106 90 169 138 164 205 115 155 196
174 116 167 169 145 116 103 74 97 90 108 67 116 116 154 220
HotpotQA 145 75 115 55 73 73 92 102 88 146 128 140 190 106 140 208
172 129 194 185 146 123 109 80 105 99 125 75 127 122 173 256
2WikiMQA 198 74 110 53 69 74 91 105 89 159 133 153 196 109 143 215
178 132 181 179 139 112 96 76 102 94 120 70 118 117 161 250
MuSiQue 108 71 108 54 72 71 89 98 85 143 124 135 189 104 138 216
178 128 206 198 148 130 113 84 112 100 126 76 134 128 173 257
GovReport 332 98 113 48 73 78 88 136 125 150 176 145 142 113 136 164
212 129 148 161 145 115 79 62 103 78 94 57 99 121 138 238
QMSum 156 85 119 51 65 63 78 116 81 146 126 136 185 105 142 218
187 121 204 201 165 134 121 85 122 105 122 70 115 121 144 207
MultiNews 459 94 115 50 67 71 88 101 99 120 126 147 168 130 156 190
202 125 147 160 134 107 78 60 101 84 95 55 102 117 146 202
TREC 214 70 112 47 69 97 93 129 98 156 142 215 220 141 213 115
143 110 166 103 105 91 90 67 84 89 139 78 129 96 192 283
TriviaQA 105 62 97 52 73 89 86 106 102 178 139 178 198 140 189 199
168 118 159 139 144 103 90 66 109 102 131 71 127 124 196 256
SAMSum 126 64 103 48 71 96 93 152 118 157 130 174 173 150 189 179
168 132 177 135 135 108 98 84 101 104 120 87 114 114 168 228
PassageCount 91 72 106 48 54 58 80 111 96 119 153 143 166 147 157 225
214 147 191 203 178 106 96 72 116 116 118 70 122 130 148 243
PassageRetrieval-en 152 92 129 70 72 75 92 121 108 128 162 149 142 139 145 175
179 145 180 172 134 113 95 87 112 109 124 91 108 118 140 238
LCC 186 59 104 44 64 80 98 128 112 159 169 161 177 173 162 163
175 133 170 147 113 102 83 69 94 80 114 84 138 130 178 247
RepoBench-P 106 63 106 46 66 89 96 132 119 168 177 169 179 183 176 189
197 138 177 148 113 102 80 71 94 80 106 83 132 133 157 221
Table 7: Per-layer token processing statistics across 32 transformer layers by DynamicKV in Mistral-7B-Instruct-v0.2 model on LongBench datasets.
A.5 Task-Aware Sensitivity Study
The presented Table 7 offers a comprehensive insight into the token processing statistics per layer for 32 transformer layers using the DynamicKV method in Mistral-7B-Instruct, as applied to various datasets from LongBench. This detailed analysis reveals significant variability in how tokens are processed at different layers depending on the specific task.
In particular, our approach demonstrates that certain tasks, such as those represented by the MultiNews and GovReport datasets, exhibit exceptionally high token processing counts even in early layers, indicating a potentially greater complexity or information density in these tasks. Conversely, datasets such as TriviaQA and SAMSum show relatively lower initial processing demands, but with significant spikes in later layers, suggesting evolving complexity throughout the document or dialogue.
Moreover, the data highlights that the DynamicKV mechanism adeptly adjusts to the unique demands of each dataset, optimizing performance and resource allocation dynamically across all layers. This adaptability is crucial for handling the diverse range of tasks and input lengths encountered in real-world applications, ensuring efficient processing without compromising on understanding or accuracy.
Algorithm 1 DynamicKV in Prefill Phase
1:  Input: initial budget K/V cache list ${K}^{b}$, ${V}^{b}$, radio max $r_{max}$, update interval $m$, mean token length $wt$, window size $ws$, sequence length $S$, head dimention $hd$, input embedding of window size ${X}^{ws}\in\mathbb{R}^{ws*d}$, initial budget Attention list computed by window token and others ${A}^{b}$,
2:  Output: Compressed K/V cache ${K}^{c}$, ${V}^{c}$
3:  $bs=(wt-ws)\times r_{max}$
4:  def Update_Buffer_Length($A$, $l$):
5:       ${A}^{gather}$leftarrow cat(([${A}$ for $l$ in (1, $l$)]), 0).view(-1)
7:       Compute the $norm$ of $cnts$, range in (0, 1)
8:       $BL$leftarrow [int(($bs*t$ / max($norm$))) for $t$ in $norm$]
9:       $r$leftarrow sum($BL$) / (($wt-ws$)$*L$)
10:       $BL$leftarrow [int($k/r$) for $k$ in $BL$]
11:       Return $BL$
12:  for $l\leftarrow 1\textbf{ to }L$ do
13:     Compute full KV states ${K}^{s}$, ${V}^{s}$
14:     for $h\leftarrow 1\textbf{ to }H$ do
16:        ${A}_{l,h}$leftarrow softmax((${X}^{ws}{W}_{h}^{Q}$) $\cdot$ ${K}_{h}^{T}$).mean(dim=-2).pooling(dim=-1)
17:     end for
18:     Append ${A}_{l}$ to ${A}^{b}$ /* current ${A}_{l}$ shape is [$H$, $S$] */
19:     /* calculate current layer buffer KV cache */
20:     indices $\leftarrow$ ${A}_{l}$.topk($bs$, dim=-1).indices.unsqueeze(-1).expand(-1, -1, $hd$)
23:     /* gradually compress*/
[…]
25:        $Bl$leftarrow Update_Buffer_Length(${A}_{l}$, $l$)
26:        /* update the buffer K/V Cache*/
27:        for $i$leftarrow 1 to $l$ do
28:           ${K}_{i}^{b}$leftarrow cat((${K}_{l}^{b}$[:,:${Bl}_{i}$,:], ${K}_{l}^{b}$[:,$-ws$:,:]), dim=-2)
29:           ${V}_{i}^{b}$leftarrow cat((${V}_{l}^{b}$[:,:${Bl}_{i}$,:], ${V}_{l}^{b}$[:,$-ws$:,:]), dim=-2)
30:        end for
31:     end if
32:  end for
33:  Update the K/V Cache ${K}^{c},{V}^{c}$ from ${K}^{b},{V}^{b}$

### v3

@@ 5 Experiments
We conduct comprehensive comparative and ablation experiments to verify the effectiveness of our DynamicKV. In § 5.1, we introduce the models, datasets and baselines used in our experiments. § 5.2 provides a performance comparison between DynamicKV and baseline approaches. Next, in § 5.3, we conduct an ablation study on the parameters of our method to validate its feasibility. We presnet the computational overhead in § 5.4. Finally, in § 5.5, we present the results of DynamicKV on the Needle in Haystack Task.
5.1 Experimental Settings
@@
LongBench is a comprehensive benchmark for evaluating the contextual understanding capabilities of LLMs. For our comparative experiments, we use 16 English datasets from this benchmark, specifically NarrativeQA (Kočiskỳ et al., 2018), Qasper (Dasigi et al., 2021), MultiFieldQA-en, HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), GovReport (Huang et al., 2021), QMSum (Zhong et al., 2021), MultiNews (Fabbri et al., 2019), TREC (Li and Roth, 2002), TriviaQA (Joshi et al., 2017), SAMSum (Gliwa et al., 2019), PassageCount, PassageRetrieval-en, LCC (Guo et al., 2023), and RepoBench-P (Liu et al., 2023).
@@
We evaluate the recent fixed-pattern token-dropping methods, including: (1) StreamingLLM Xiao et al. (2023), which utilizes attention sinks and rolling KV caches to retain the most recent tokens. (2) H2O Zhang et al. (2024b), which employs a Heavy Hitter Oracle for KV cache eviction. (3) SnapKV Li et al. (2024), which selects important tokens for each attention head through clustering. (4) PyramidKV Zhang et al. (2024a), which introduces a pyramid pattern where layers select important tokens in a monotonically decreasing manner.
@@ 5.2 Comparative Experiments on LongBench
Model Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
@@
Table 1: Performance comparison on the LongBench dataset for full KV cache, previous methods (StreamingLLM, H2O, SnapKV, PyramidKV), and our DynamicKV method, with KV cache sizes of 512, using models including LLaMA3-8B-Instruct, Mistral-7B-Instruct-v0.2, QWen2-7B-Instruct, and InternLM-2.5-Chat-1M. Bold indicates the best performance.
With the total KV cache size constrained to just 512, we evaluate the performance retention of StreamingLLM, H2O, SnapKV, PyramidKV, and our proposed approach, DynamicKV, relative to the FullKV. As shown in Table 1, DynamicKV consistently outperforms existing methods, enven when operating with an exceptionally low cache-to-context ratio of only 6.9%. Notably, DynamicKV exceeds the best-performing baseline by 0.43%, 0.19%, 0.69%, and 0.53% across comparable models – retaining 97%, 96%, 96%, and 89% of FullKV’s performance, respectively. These results underscore DynamicKV’s remarkable ability to preserve near FullKV-level performance under extreme memory constraints. Further more, DynamicKV not only matches but enhances PyramidKV’s capabilities on complex tasks such as code completion, significantly extending the performance ceiling at lower cache capacities. In addition, we also compared the performance with a KV cache size of 128. The detailed results can be found in Appendix A.5.
5.3 Ablation Study
In this study, we investigate the performance of the DynamicKV mechanism across varying key-value cache sizes. The results, as shown in Figure 4, reveal a consistent improvement in performance with an increase in the cache size for all evaluated models. For the LlaMA-3-8B-Instruct, the performance metric improved from 34.93 to 41.22 as the key-value cache size was increased from 64 to 1024. This improvement is also applicable to other models. These findings underscore the effectiveness of the DynamicKV cache in leveraging KV cache compression to maintain the capabilities of long context. Notably, a larger cache capacity is generally associated with superior performance. Nonetheless, it is essential to strike a balance when selecting the cache size, taking into account the practical constraints related to storage and computational resources.
Figure 4: Performance of DynamicKV with different KV cache size on LongBench. The evaluation metrics are the average score of LongBench across datasets.
5.4 Computational Overhead
To better understand the overhead of our DynamicKV, we compare the computational overhead with the FullKV using Llama on LongBench. The evaluation metrics are Time-to-First-Token (TTFT), Time-Per-Output-Token (TPOT), end-to-end latency, and GPU memory usage (GB). We present the result in Table 2.
We can observe that DynamicKV deliver 129% higher TPOT, 56% lower latency comparison with FullKV. Experimental results show that our DynamicKV offers significant advantages in both computational efficiency and memory usage. More efficient experimental results can be found in Appendix A.4.
[…]
FullKV 3.52 11.65 706.56 30.48
DynamicKV 3.58 26.69 310.56 27.06
Table 2: Efficiency comparison between FullKV and DynamicKV. We conduct experiments with a fixed context window ($m=128$), the input length is $32$K and output length is $8$K.
@@ 5.5 Visualization on Needle-in-Haystack Task
Figure 5: Performance Comparison on the Needle in a Haystack Task using Mistral-7B-Instruct-v0.2 with 32k context size in 64 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the length.
We evaluate the in-context retrieval capabilities of LLMs using the “Fact Retrieval Across Context Lengths” benchmark (also known as Needle In A Haystack) – a challenging dataset designed to assess whether a model can accurately extract key information from long input sequences. To this end, we adopt Mistral as the base model and extend the context length up to 32K tokens. We compare multiple KV cache compression strategies, including StreamingLLM, PyramidKV, and our proposed DynamicKV, at cache sized of 64 and the FullKV baseline. The results, shown in Figure 5, highlight that DynamicKV retains 90% of the model’s original performance even under aggressive compression – achieving accuracy gains of 57%, 37%, 41% and 11% over competing methods.
Moreover, the results demonstrate that at context lengths up to 7K tokens, DynamicKV’s extreme compression nearly achieves full accuracy. Beyond this range, it continues to significantly outperform all baselines. These results underscore DynamicKV’s superior capability in hierarchical token selection, and validate our hypothesis that the distribution of critical tokens across layers is inherently dynamic.
☞ A Note on More Details in the Appendix
See Appendix A.1 and A.2 for a more detailed description of the experimental settings, Appendix A.3 for additional results from Need in a HayStack, Appendix A.4 for efficiency experiments and Appendix A.5 for result of KV cache size of 128 on the LongBench dataset.
@@ 6 Conclusion
We investigate task-specific attention patterns in LLMs processing long-context inputs and find distinct attention distributions across tasks. To address this, we propose DynamicKV, a layer-adaptive KV cache compression framework that dynamically optimizes KV cache allocation per layer. We evaluate the effectiveness and generalizability of DynamicKV through experiments on 16 datasets from the LongBench benchmark, demonstrating its broad applicability and performance benefits. From the results, we mainly conclude that: (1) a wave-like pattern is followed in complex reasoning tasks (e.g., code completion tasks); (2) a pyramid-like pattern is followed in Synthetic and Summarization tasks; (3) The dynamic hierarchical adaptive DynamicKV approach is capable of formulating a relatively appropriate KV cache retention strategy in accordance with diverse tasks. Particularly, in the circumstance of maintaining an extremely small KV cache size, the effect is significantly enhanced. In the future, we hope that there is a more suitable method to perform KV cache compression without increasing the computation.
@@
Ethics and Reproducibility Statements
Ethics
We take ethical considerations seriously and follow the guidelines outlined by the ACL Ethics Policy. The DynamicKV method is designed to optimize long-context inference in LLMs, without the need for collecting sensitive or private information. All datasets used in the experiments are publicly available and widely adopted by the research community, ensuring transparency and accessibility. We do not foresee any significant ethical concerns related to the development and use of the DynamicKV method.
Reproducibility
To ensure reproducibility, we provide detailed descriptions of our experimental setup, including model configurations, datasets, and performance metrics. Furthermore, we have provided our code in the Supplementary Material. We hope that the provided resources will support further advancements in efficient LLM inference and memory management.
@@ Appendix A Appendix
This appendix presents a detailed description of the used models and dataset (Appendix A.1 and A.2), along with additional results from Need in a HayStack (Appendix A.3), comprehensive efficiency experiments (Appendix A.4), and more experimenet results on LongBench (Appendix A.5).
@@ A.1 Model Details
Our experiments are based on four representative open-sourced LLMs, namely LlaMA-3-8B-Instruct, Mistral-7B-Instruct-v0.2, Qwen2-7B-Instruct, and InternLM2.5-Chat-1M. Testing examples are evaluated in a generative format, with answers generated by greedy decoding across all tasks to ensure a fair comparison. All the model structures and details in our experiment are shown in Table 3.
@@ A.2 Dataset Details
We evaluate the performance of DynamicKV on long-context tasks using LongBench Bai et al. (2023), a rigorously constructed benchmark suite designed to challenge language models with extended documents and intricate information sequences. Developed for comprehensive, multi-task assessment, LongBench serves as a critical tool for measuring a model’s ability to understand and reason over long-context inputs with precision and depth. The data sources, average length, evaluation metrics, language, and data volume of subdatasets of LongBench are shown in Table 4.
@@
Table 5: Comparison of different KV cache compression methods in the Needle in a Haystack task.
@@ A.3 Need in a HayStack
As shown in Table 5, we compare the performance of various KV cache compression methods – StreamingLLM, H2O, SnapKV, PyramidKV, and DynamicKV – on the Needle in a Haystack task using two models: LlaMA-3-8B-Instruct and Qwen-2-7B-Instruct. Across both models, our DynamicKV achieves the highest performance, scoring 0.9 for LlaMA-3-8B-Instruct and 0.87 for Qwen-2-7B-Instruct. These results highlight DynamicKV’s superior ability to retain task-critical information in long-context scenarios.
@@ A.4 Efficiency Experiments
A.5 More Experiment Result on LongBench
Model Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
StreamingLLM 17.85 9.50 23.09 37.84 29.02 16.77 17.91 20.42 20.16 44.00 73.00 30.00 5.80 69.50 48.38 49.31 32.03
StreamingLLM 16.58 14.76 30.36 28.13 21.76 11.98 18.26 19.02 19.16 43.50 74.12 28.50 2.50 31.81 43.65 41.19 27.83
StreamingLLM 19.25 23.63 26.51 14.00 15.30 7.46 18.07 19.30 18.30 47.00 77.92 31.57 6.50 17.00 42.52 41.94 26.64
StreamingLLM 17.91 13.02 24.31 24.27 16.01 11.29 17.29 20.62 18.06 48.5 67.53 21.93 0.82 87.39 43.45 42.79 29.70
Table 7: Performance comparison on the LongBench dataset for full KV cache, previous methods (StreamingLLM, H2O, SnapKV, PyramidKV), and our DynamicKV method, with KV cache sizes of 128, using models including LLaMA3-8B-Instruct, Mistral-7B-Instruct-v0.2, QWen2-7B-Instruct, and InternLM-2.5-Chat-1M. Bold indicates the best performance.
Table 7 presents a performance comparison on the LongBench for different KV cache compression methods (StreamingLLM, H2O, SnapKV, PyramidKV and our DynamicKV) with a fixed cache size of 128. We conduct experiments across various tasks such as Single-Document QA, Multi-Document QA, Summarization, Few-shot Learning, Synthetic tasks, and Code Completion.
The results show that our DynamicKV consistently achieves competitive or superior performance compared to previous methods. While FullKV yields the highest average scores, DynamicKV achieves the best or near-best performance across several models – particularly excelling with Mistral-7B-Instruct-v0.2 and InternLM-2.5-Chat-1M – demonstrating effective memory compression with minimal loss in accuracy.